# Data Mixing Agent: Learning to Re-weight Domains for Continual Pre-training

### v1

@@ 4.1 Experimental Settings
For data from the source field, the self-pretrained LLaMA-3B models utilize their corresponding pre-training data, each with 100B tokens. We use randomly sampled data from the Pythia-1.4B model as the source field data for itself, applying the agent in scenarios where the data from the source field is not directly available. Following the method in Sec. 3.1, we sample 10B tokens by pre-pending the start-of-sentence token to start generation with a default temperature of 1.0. The generated data are then filtered through the Nvidia text quality classifier  [^4] https://huggingface.co/nvidia/quality-classifier-deberta, where all data within the "Low quality" class are discarded, resulting in a source field dataset with around 7.7B tokens. For data from the math reasoning field, we select the math split (10B tokens) of the Dolmino-mix-1124 dataset  [^5] https://huggingface.co/datasets/allenai/dolmino-mix-1124, which was used for the mid-training process of the OLMo2 model series (OLMo et al., 2024), including data sources such as TuluMath (Ivison et al., 2024), MathCoder (Wang et al., 2023), and Metamath (Yu et al., 2023). For data from the code generation field, we select the GitHub training split  [^6] https://huggingface.co/datasets/MBZUAI-LLM/SlimPajama-627B-DC of the SlimPajama-DC dataset (Shen et al., 2023) with 30B tokens.
Evaluation Benchmarks
@@
We evaluate target models’ general capabilities by evaluating with the lm__\__eval evaluation library  [^7] https://github.com/EleutherAI/lm-evaluation-harness on the MMLU (Hendrycks et al., 2020), HellaSwag (Zellers et al., 2019) (Hella.), OpenBookQA (Mihaylov et al., 2018) (OBQA), Winogrande (Sakaguchi et al., 2021) (Wino.), ARC-Challenge (Clark et al., 2018) (ARC-C), PiQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), and LogiQA (Liu et al., 2020) benchmarks. We evaluate the math reasoning capabilities using the math__\__lm__\__eval library  [^8] https://github.com/ZubinGou/math-evaluation-harness on the GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), Minerva (Lewkowycz et al., 2022), and MathQA (Amini et al., 2019) benchmarks. We evaluate the code generation capabilities using the eval__\__plus library  [^9] https://github.com/evalplus/evalplus on the HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) benchmarks. The MMLU and GSM8K benchmarks are evaluated with a 5-shot setting, the Minerva benchmark is evaluated in a 4-shot setting. Other benchmarks are evaluated with a zero-shot setting.
Target Model Training Setting
@@
Firstly, the agent is trained based on a 26-dimensional domain definition, leading to a 52-dimensional domain reweighting space (shown in Sec. 3.1). To evaluate on different domain spaces, we further employ the data mixing agent on the original 2-dimensional space based on data sources (source and target). Note that this action does not require the agent to retrain, as its action can be directly converted by summing the 26 probabilities for source/target fields into a single dimension, still preserving a probability distribution. During training, we set the number of reweighting samples per step $R_{tgt}$ to 64K and the maximum data reweighting steps $M_{tgt}$ to 80. We use the same evaluation environment as the Data Mixing Agent training when the target field is math reasoning, and change the MATH validation set to 1,000 random samples from the GitHub validation split of the SlimPajama-DC dataset. Due to resource limits, we only train on the 2-dimensional data reweighting space for code generation.
Implementation
@@
4.2 Evaluation Results on Math Reasoning
The evaluation results of continual pretraining on the math reasoning target field are shown in Table 1. According to the results, we have the following observations:
@@ 4.2 Evaluation Results on Math Reasoning
Naive training significantly improves target model performance on the target field but leads to drastic collapse on the capabilities of the source field. Compared to the base model, the average math reasoning performance increases by an average of 22.77% on the four target models, indicating the effectiveness of training on high-quality in-distribution data for the target field. However, the performance on general benchmarks drops by an average of 11.96%, showing a significant degradation in the source-field model capability. These results further highlight the existence of catastrophic forgetting problems in continual pre-training scenarios, motivating exploration in data mixture and domain reweighting algorithms.
@@
Domain reweighting algorithms such as RegMix can achieve balanced performance across fields. According to the results, the RegMix method exhibits a trade-off effect across domains. On the 2-dimensional data reweighting space, it outperforms the base model on math reasoning by an average of 18.47%, while largely preserving general capabilities with a mere 2.28% degradation on the corresponding benchmarks. RegMix also outperforms the naive training by 5.03% on the overall average performance. Similar conclusions can be drawn from the results on the 52-dimensional domain space in Fig.1(b). These results show that the catastrophic forgetting problem can be considerably alleviated by carefully curating data mixtures of source and target fields.
@@
Data Mixing Agent significantly outperforms other methods in balanced performance across fields. In Fig. 1(a), for the in-distribution LLaMA-3B-DCLM-100B target model, DataAgentRL outperforms the RegMix results on 7 out of 8 general benchmarks and all 4 math benchmarks. It achieves the best average performance 54.04% and 33.02% on general/math benchmarks, even outperforming the base model in general ability and the naively trained model on math reasoning. These results prove that DataAgentRL can effectively curate the data mixture to improve both general and math reasoning capabilities. With careful domain reweighting, increasing capability on the target field can further enhance performance on the source field. Overall, DataAgentRL achieves 47.03% on average, surpassing RegMix by 3.02% and the base model by 8.88%. DataAgentRL also outperforms DataAgentSFT by a large margin of 2.08%. This advantage shows that the empirical guidance presented in Algorithm 1 is trivial compared to heuristics derived from the broader sampling of data mixing trajectories, proving reinforcement learning with CQL as a crucial step towards capable agents.
@@
The capabilities of data mixing agents can generalize across target models, source-field data, and domain spaces without retraining. Though our data mixing agent is trained on the 52-dimensional data reweighting space with trajectories sampled with the DCLM data, it effectively guides domain reweighting for four target models across 2 domain definitions. For example, in Fig. 1(a), DataAgentRL outperforms RegMix by an average of 1.66% on the two unseen target models: LLaMA-3B-FWE-100B and LLaMA-3B-Nemotron-100B, based on the 2-dimensional domain space. In Fig. 1(b), DataAgentRL outperforms RegMix by an average of 1.41% based on the 52-dimensional domain space. These results indicate that Data Mixing Agent learns data- and model-agnostic heuristics from the sampled trajectories that can guide domain reweighting on multiple source-field data distributions, which is crucial to the efficiency of this algorithm, as the feedback collection for sampled data trajectories (Sec. 3.1) requires considerable computations. With these generalization capabilities, the agent is still expected to perform well in applications to new target models and source-field data without re-training.
@@
Data mixing agent is effective in guiding continual pre-training on estimated start state and data mixtures with synthetic source-field data. We prove this by reweighting domains on the Pythia-1.4B target model with the estimated start state obtained as in Sec. 3.1 and source-field data obtained as described in Sec. 4.1. In math reasoning, DataAgentRL also outperforms RegMix by 0.59% in Table 1(a) and 3.7% in Table 1(b). However, the preservation on general capabilities significantly drops, with a 4.69% and 5.65% gap on 2-dimensional and 52-dimensional domain spaces. Overall, DataAgentRL still significantly improves average performance compared to the base model, and outperforms RegMix by 1.05% in Table 1(a) and 0.61% in Table 1(b).
4.3 Evaluation Results on Code Generation
@@
The capabilities of Data Mixing Agent can partially generalize across target fields without retraining. DataAgentRL achieves the best average performance of 46.3% and 41.63% on the LLaMA-3B-DCLM-100B and Pythia-1.4B target models, outperforming the RegMix method by 1.45% and 2.67%. These results prove that heuristics learned in the math reasoning field can be partially transferred to the code generation field without modifying the weights of the agent. However, we observe a degradation in DataAgentRL’s advantage over the baseline methods in code generation. For example, DataAgentRL outperforms naive training by 6.22%, while in Table 1(a), the advantage is 8.52%. This is mainly due to that applying DataAgentRL to code generation leads to a major 3.18% drop on general benchmarks compared to math reasoning, which indicates the existence of heuristics that are dependent on the target field and the potential misalignment when converting them to a new target field.
@@
Figure 4: The two data mixing agents’ output domain reweighting trajectories based on the 2-dimensional domain space, training on the LLaMA-3B-DCLM-100B model and the math reasoning field. The dashed line denotes the optimal domain distributions determined by RegMix.
Figure 5: DataAgentSFT’s domain reweighting trajectories based on the 52-dimensional domain space, training on the LLaMA-3B-DCLM-100B model and the math reasoning field. The legends within each sub-figure are the same as those of Fig. 4.
Figure 6: DataAgentRL’s domain reweighting trajectories based on the 52-dimensional domain space, training on the LLaMA-3B-DCLM-100B model and the math reasoning field. The legends within each sub-figure are the same as those of Fig. 4.
4.4 Analysis on Domain Reweighting Trajectories
In this section, we showcase the domain reweighting process guided by Data Mixing Agent to train the LLaMA-3B-DCLM-100B model on the math reasoning field, aiming to provide more intuitions on its actions based on the heuristics and feedback. The trajectories on the 2-dimensional and 52-dimensional domain spaces are provided in Fig. 4, Fig. 5, and Fig. 6. We have the following observations:
@@ 4.4 Analysis on Domain Reweighting Trajectories
Data Mixing Agents learn heuristics and perform actions that correspond to human intuitions on the target capabilities. Our work uses the MMLU evaluation set to represent the general capabilities in the environment. Wettig et al. (2025) summarized the top-3 domains that benefit the MMLU performance: $Science\&Tech.$, $Health$, and $Politics$. In Fig. 6, we observe a significant uplift of the target data distributions in the corresponding domains compared to the RegMix domain distributions: $Science$, $Health$, and $People\&Society$. Wettig et al. (2025) also enumerated domains that can hurt performance on MMLU, such as $Fashion\&Beauty$, while DataAgentRL also conveys an explicit down-sampling process in the $Beauty\&Fitness$ domain. These observations further ensure the effectiveness of the learned heuristics, encouraging the discovery of more heuristics via the agent’s trajectories. For example, DataAgentRL continuously reduces data from both source and target fields in the $Pets\&Animals$ domain, possibly indicating its lack of importance in enhancing either general or math reasoning capabilities.
@@
4.5 Data Efficiency
We explore how efficiently the Data Mixing Agents leverage the source and target field data to improve or preserve model capabilities in the corresponding fields. Training on the mixture of DCLM-100B and the math split of Dolmino-mix-1124 datasets, we record the performance dynamics of the LLaMA-3B-DCLM-100B target model on the general/math evaluation environment with increasing training data (measured in Billion tokens) on the general/math reasoning field. The results are shown in Fig. 7. We have the following observations:
@@ 4.5 Data Efficiency
5 Related Work
5.1 Continual Pre-training

### v2

@@ D.3 Target Model and Data
For data from the source field, the self-pretrained LLaMA-3B models utilize their corresponding pre-training data, each with 100B tokens. We use randomly sampled data from the Pythia-1.4B model as the source field data for itself, applying the agent in scenarios where the data from the source field is not directly available. Following the method in Appendix B, we sample 10B tokens by pre-pending the start-of-sentence token to start generation with a default temperature of 1.0. The generated data are then filtered through the Nvidia text quality classifier  [^5] https://huggingface.co/nvidia/quality-classifier-deberta, where all data within the "Low quality" class are discarded, resulting in a source field dataset with around 7.7B tokens. For data from the math reasoning field, we select the math split (10B tokens) of the Dolmino-mix-1124 dataset  [^6] https://huggingface.co/datasets/allenai/dolmino-mix-1124, which was used for the mid-training process of the OLMo2 model series (OLMo et al., 2024), including data sources such as TuluMath (Ivison et al., 2024), MathCoder (Wang et al., 2023), and Metamath (Yu et al., 2023). For data from the code generation field, we select the GitHub training split  [^7] https://huggingface.co/datasets/MBZUAI-LLM/SlimPajama-627B-DC of the SlimPajama-DC dataset (Shen et al., 2023) with 30B tokens.
D.4 Elaborations on Target Model Selection
Our experiments aim to validate the proposed data mixing agent as a domain reweighting method rather than to achieve state-of-the-art leaderboard performance. Therefore, absolute performance gaps to SOTA LLMs do not undermine the conclusions. Specifically, we employ 3B-parameter models pretrained on 100B tokens, whereas strong open-source LLMs (e.g., Llama-3.1-405B-Instruct) use hundreds of billions of parameters and are pretrained on trillions of tokens. We avoid larger open-source LLMs for three reasons: (1) many LLMs have been heavily optimized for math or code, introducing biases that confound controlled evaluation of domain reweighting; (2) their pretraining data compositions are not publicly available, making them unsuitable for continual pretraining studies; and (3) their large scale (typically>70B parameters) renders continual pretraining prohibitively expensive.
D.5 Evaluation Benchmarks
@@
We evaluate target models’ general capabilities by evaluating with the lm_\_eval evaluation library  [^8] https://github.com/EleutherAI/lm-evaluation-harness on the MMLU (Hendrycks et al., 2020), HellaSwag (Zellers et al., 2019) (Hella.), OpenBookQA (Mihaylov et al., 2018) (OBQA), Winogrande (Sakaguchi et al., 2021) (Wino.), ARC-Challenge (Clark et al., 2018) (ARC-C), PiQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), and LogiQA (Liu et al., 2020) benchmarks. We evaluate the math reasoning capabilities using the math_\_lm_\_eval library  [^9] https://github.com/ZubinGou/math-evaluation-harness on the GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), Minerva (Lewkowycz et al., 2022), and MathQA (Amini et al., 2019) benchmarks. We evaluate the code generation capabilities using the eval_\_plus library  [^10] https://github.com/evalplus/evalplus on the HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) benchmarks. The MMLU and GSM8K benchmarks are evaluated with a 5-shot setting, the Minerva benchmark is evaluated in a 4-shot setting. Other benchmarks are evaluated with a zero-shot setting.
D.6 Target Model Training Setting
@@
Firstly, the agent is trained based on a 26-dimensional domain definition, leading to a 52-dimensional domain reweighting space. To evaluate on different domain spaces, we further employ the data mixing agent on the original 2-dimensional space based on data sources (source and target). Note that this action does not require the agent to retrain, as its action can be directly converted by summing the 26 probabilities for source/target fields into a single dimension, still preserving a probability distribution. During training, we set the number of reweighting samples per step $R_{tgt}$ to 64K and the maximum data reweighting steps $M_{tgt}$ to 80. We use the same evaluation environment as the Data Mixing Agent training when the target field is math reasoning, and change the MATH validation set to 1,000 random samples from the GitHub validation split of the SlimPajama-DC dataset when the target field is code generation. Due to resource limits and the highly imbalanced distribution of code data in the 26-dimensional domain space, we only train on the 2-dimensional data reweighting space for code generation.
D.7 Implementation
@@ Appendix E Experimental Results
E.1 Full Results and Analysis on Math Reasoning
The evaluation results of continual pretraining on the math reasoning target field are shown in Table 3. The data agent and all domain reweighting baseline methods take significantly more GPU training hours than naive training. The main reason is that these domain reweighting methods require a mixture of target-field data and source-field data, while naive training only requires target-field data. By training on more tokens, domain reweighting methods significantly mitigate the catastrophic forgetting problem encountered in naive training, usually outperforming naive training by a large margin on average performance. According to the results, we have the following observations:
@@ E.1 Full Results and Analysis on Math Reasoning
Naive training significantly improves target model performance on the target field but leads to drastic collapse on the capabilities of the source field. Compared to the base model, the average math reasoning performance increases by an average of 22.77% on the four target models, indicating the effectiveness of training on high-quality in-distribution data for the target field. However, the performance on general benchmarks drops by an average of 11.96%, showing a significant degradation in the source-field model capability. These results further highlight the existence of catastrophic forgetting problems in continual pre-training scenarios, motivating exploration in data mixture and domain reweighting algorithms. While data mixing agent bears an average of 0.57% gap to naive training in the target domain, it outperforms naive training by an average of 11.99% in the source domain, showing a significant mitigation of the catastrophic forgetting problem.
@@
Domain reweighting algorithms such as RegMix can achieve balanced performance across fields. According to the results, the RegMix method exhibits a trade-off effect across domains. On the 2-dimensional data reweighting space, it outperforms the base model on math reasoning by an average of 18.47%, while largely preserving general capabilities with a mere 2.28% degradation on the corresponding benchmarks. RegMix also outperforms the naive training by 5.03% on the overall average performance. Similar conclusions can be drawn from the results on the 52-dimensional domain space in Table 3(b). These results show that the catastrophic forgetting problem can be considerably alleviated by carefully curating data mixtures of source and target fields.
@@
Data Mixing Agent significantly outperforms other methods in balanced performance across fields. In Table 3(a), for the in-distribution LLaMA-DCLM target model, DataAgentRL outperforms the RegMix results on 7 out of 8 general benchmarks and all 4 math benchmarks. It achieves the best average performance 54.04% and 33.02% on general/math benchmarks, even outperforming the base model in general ability and the naively trained model on math reasoning. These results prove that DataAgentRL can effectively curate the data mixture to improve both general and math reasoning capabilities. With careful domain reweighting, increasing capability on the target field can further enhance performance on the source field. Overall, DataAgentRL achieves 47.03% on average, surpassing RegMix by 3.02%, DBL by 3.53%, and the base model by 8.88%. DataAgentRL also outperforms DataAgentSFT by a large margin of 2.08%. This advantage shows that the empirical guidance presented in the trajectory sampling algorithm is trivial compared to heuristics derived from the broader sampling of data mixing trajectories, underscoring the importance of reinforcement learning with CQL as a crucial step towards capable agents.
@@
The capabilities of data mixing agents can generalize across target models, source-field data, and domain spaces without retraining. Though our data mixing agent is trained on the 52-dimensional data reweighting space with trajectories sampled with the DCLM data, it effectively guides domain reweighting for four target models across 2 domain definitions. For example, in Table 3(a), DataAgentRL outperforms RegMix by an average of 1.66%, DBL by an average of 2.37% on the two unseen target models: LLaMA-FWE and LLaMA-Nemo, based on the 2-dimensional domain space. In Table 3(b), DataAgentRL outperforms RegMix by an average of 1.41% based on the 52-dimensional domain space. These results indicate that Data Mixing Agent learns data- and model-agnostic heuristics from the sampled trajectories that can guide domain reweighting on multiple source-field data distributions, which is crucial to the efficiency of this algorithm, as the feedback collection for sampled data trajectories requires considerable computations. With these generalization capabilities, the agent is still expected to perform well in applications to new target models and source-field data without re-training.
@@
Data mixing agent is effective in guiding continual pre-training on estimated start state and data mixtures with synthetic source-field data. We prove this by reweighting domains on the Pythia-1.4B target model with the estimated start state obtained as in Appendix B and source-field data obtained as described in Appendix D.3. In math reasoning, DataAgentRL also outperforms RegMix by 0.59% in Table 3(a) and 3.7% in Table 3(b). However, the preservation on general capabilities significantly drops, with a 4.69% and 5.65% gap on 2-dimensional and 52-dimensional domain spaces. Overall, DataAgentRL still significantly improves average performance compared to the base model, and outperforms RegMix by 1.05% in Table 3(a) and 0.61% in Table 3(b).
E.2 Full Results and Analysis on Code Generation
@@
The capabilities of Data Mixing Agent can partially generalize across target fields without retraining. DataAgentRL achieves the best average performance of 46.3% and 41.63% on the LLaMA-DCLM and Pythia-1.4B target models, outperforming the DBL method by an average of 2.78% and RegMix method by 2.67%. These results prove that heuristics learned in the math reasoning field can be partially transferred to the code generation field without modifying the weights of the agent. However, we observe a degradation in DataAgentRL’s advantage over the baseline methods in code generation. For example, DataAgentRL outperforms naive training by 6.22%, while in Table 3(a), the advantage is 8.52%. This is mainly due to that applying DataAgentRL to code generation leads to a major 3.18% drop on general benchmarks compared to math reasoning, which indicates the existence of heuristics that are dependent on the target field and the potential misalignment when converting them to a new target field.
@@
Figure 5: The two data mixing agents’ output domain reweighting trajectories based on the 2-dimensional domain space, training on the LLaMA-DCLM model and the math reasoning field. The dashed line denotes the optimal domain distributions determined by RegMix.
Figure 6: DataAgentSFT’s domain reweighting trajectories based on the 52-dimensional domain space, training on the LLaMA-DCLM model and the math reasoning field. The legends within each sub-figure are the same as those of Fig. 5.
Figure 7: DataAgentRL’s domain reweighting trajectories based on the 52-dimensional domain space, training on the LLaMA-DCLM model and the math reasoning field. The legends within each sub-figure are the same as those of Fig. 5.
E.3 Full Analysis on Domain Reweighting Trajectories
In this section, we showcase the domain reweighting process guided by Data Mixing Agent to train the LLaMA-DCLM model on the math reasoning field, aiming to provide more intuitions on its actions based on the heuristics and feedback. The trajectories on the 2-dimensional and 52-dimensional domain spaces are provided in Fig. 5, Fig. 6, and Fig. 7. We have the following observations:
@@ E.3 Full Analysis on Domain Reweighting Trajectories
Data Mixing Agents learn heuristics and perform actions that correspond to human intuitions on the target capabilities. Our work uses the MMLU evaluation set to represent the general capabilities in the environment. Wettig et al. (2025) summarized the top-3 domains that benefit the MMLU performance: Science&Tech.Science\&Tech., HealthHealth, and $Politics$. In Fig. 7, we observe a significant uplift of the target data distributions in the corresponding domains compared to the RegMix domain distributions: $Science$, HealthHealth, and People&SocietyPeople\&Society. Wettig et al. (2025) also enumerated domains that can hurt performance on MMLU, such as Fashion&BeautyFashion\&Beauty, while DataAgentRL also conveys an explicit down-sampling process in the Beauty&FitnessBeauty\&Fitness domain. These observations further ensure the effectiveness of the learned heuristics, encouraging the discovery of more heuristics via the agent’s trajectories. For example, DataAgentRL continuously reduces data from both source and target fields in the Pets&AnimalsPets\&Animals domain, possibly indicating its lack of importance in enhancing either general or math reasoning capabilities.
@@
E.4 Full Analysis on Data Efficiency
We explore how efficiently the Data Mixing Agents leverage the source and target field data to improve or preserve model capabilities in the corresponding fields. Training on the mixture of DCLM-100B and the math split of Dolmino-mix-1124 datasets, we record the performance dynamics of the LLaMA-DCLM target model on the general/math evaluation environment with increasing training data (measured in Billion tokens) on the general/math reasoning field. The results are shown in Fig. 8. We have the following observations:
@@ E.4 Full Analysis on Data Efficiency
E.5 Consistency between Rewards and Model Performance
To examine whether the proposed environment feedback provides a consistent and reliable signal for downstream model performance, we analyze multiple checkpoints along the LLaMA-DCLM training trajectory in the math domain. Specifically, we evaluate five checkpoints on different training stages (with 3.8B to 19.92B trained tokens) and report their corresponding environment feedback signals (denoted as $Env_{general}$ and $Env_{math}$) and average benchmark performance on general-domain and math tasks (denoted as bench_general_avgbench\_general\_avg and bench_math_avgbench\_math\_avg).
Training Tokens $Env_{general}$ $Env_{math}$ bench_general_avg bench_math_avg
3.8B -2.558 -1.486 52.96 10.02
7.64B -2.564 -1.058 51.08 20.70
12.84B -2.591 -0.896 48.00 32.95
16.52B -2.473 -0.950 58.22 23.10
19.92B -2.500 -0.902 54.04 33.02
Table 5: Environment feedback signals and benchmark performance at different training stages in the math domain, along the LLaMA-DCLM trajectory.
The results are shown in Table 5. Across these checkpoints, we observe a strong monotonic relationship between the environment feedback and final benchmark performance. The Pearson’s correlation coefficient between $Env_{general}$ and bench_general_avg is 0.943, while the correlation between $Env_{math}$ and bench_math_avg reaches 0.919. These high correlations indicate that the environment feedback closely tracks the model’s eventual evaluation accuracy, validating its consistency as a training signal. This result demonstrates that the reward used by the data mixing agent is well aligned with downstream task performance, thereby supporting the reliability and effectiveness of our method.
Appendix F Related Work
F.1 Continual Pre-training